OpenStudy (anonymous):

please lend a hand for the last 5 problems on my last post would greatly appreciate it

OpenStudy (anonymous):

\[\frac{\sqrt[3]{56}}{\sqrt[3]{7}}=\sqrt[3]{8}\]\[\sqrt[3]{8} =\sqrt[3]{2^3}=2\]